### v1

# Neuromorphic Intermediate Representation: A Unified Instruction Set for Interoperable Brain-Inspired Computing

Abstract
Spiking neural networks and neuromorphic hardware platforms that emulate neural dynamics are slowly gaining momentum and entering main-stream usage. Despite a well-established mathematical foundation for neural dynamics, the implementation details vary greatly across different platforms. Correspondingly, there are a plethora of software and hardware implementations with their own unique technology stacks. Consequently, neuromorphic systems typically diverge from the expected computational model, which challenges the reproducibility and reliability across platforms. Additionally, most neuromorphic hardware is limited by its access via a single software frameworks with a limited set of training procedures. Here, we establish a common reference-frame for computations in neuromorphic systems, dubbed the Neuromorphic Intermediate Representation (NIR). NIR defines a set of computational primitives as idealized continuous-time hybrid systems that can be composed into graphs and mapped to and from various neuromorphic technology stacks. By abstracting away assumptions around discretization and hardware constraints, NIR faithfully captures the fundamental computation, while simultaneously exposing the exact differences between the evaluated implementation and the idealized mathematical formalism. We reproduce three NIR graphs across 7 neuromorphic simulators and 4 hardware platforms, demonstrating support for an unprecedented number of neuromorphic systems. With NIR, we decouple the evolution of neuromorphic hardware and software, ultimately increasing the interoperability between platforms and improving accessibility to neuromorphic technologies. We believe that NIR is an important step towards the continued study of brain-inspired hardware and bottom-up approaches aimed at an improved understanding of the computational underpinnings of nervous systems. NIR is available at https://github.com/neuromorphs/NIR
1 Introduction
The human brain is an exemplar of computational efficiency, outperforming today’s advanced machine learning models and hardware in numerous measures, particularly regarding energy consumption. State-of-the-art machine learning algorithms require vast amounts of energy and computational resources, especially when trained on large datasets, whereas the brain operates on a tiny power budget, performing a myriad of complex tasks seamlessly. Unlike digital computations that rely on precise binary logic, the brain’s computations are predominantly analog in nature, utilizing gradients of ion concentrations and single-bit activations, commonly referred to as spikes. The biological circuits in the brain can be described as dynamical systems that evolve over time and are inherently noisy, yet they manage to process information with remarkable accuracy and adaptability [1, 2, 3]. Neuromorphic computing, defined as “circuits that emulate the temporal processing of signals in the brain” [2, p. 361], draws inspiration from the efficiency of the nervous systems to rethink the principles of information processing. As such, neuromorphic computing stands in contrast to the conventional and widespread von Neumann architecture, relying on entirely new algorithms, software tools and hardware [4].
Figure 1: High-level overview of what NIR is. (a) NIR allows for continuous-time representation of nodes that can then be executed on continuous-time hardware or simulators, or discretized for use on discrete-time hardware or simulators. (b) A taxonomy of discrete and continuous time hardware and simulators.
Neuromorphic systems are typically approached bottom-up or top-down [5]. Rapid scientific progress in both directions means that neuromorphic platforms need to compete on both efficient and flexibility. Secondly, contemporary neuromorphic hardware and software are often developed together, similar to the biological system from which we draw inspiration—evidently a product of conjoined development between physical form and behavioral function [2]. blurring the lines between top-down function and bottom-up computation. The blurring of top-down and bottom-up approaches makes it difficult to extrapolate clear abstractions that generalize to other technology stacks, which impedes incremental technological progress. As a result, we face a landscape of heterogeneous neuromorphic simulators and design tools. By identifying domain-specific computational primitives, we argue that a layer of abstraction can be established to achieve greater interoperability and, in turn: (1) allow hardware and software efforts to develop independently and more rapidly, (2) lower the barrier-of-entry for newcomers, and (3) provide theoreticians with a more stable representation to reason about neuromorphic circuits.
The field of deep learning faced similar incompatibility challenges a decade ago, but since then, intermediate representations and compiler frameworks such as ONNX [6], MLIR [7], XLA [8], and TVM [9] have bridged the gap between model description and different hardware accelerators. Specifically, these representations map models written in higher-level languages to different classes of hardware backends, such as CPUs, GPUs, TPUs, and FPGAs. Given the nature of clocked computation in the traditional computing stack, deep learning representations are based on digital instructions in discrete time. This contrasts our definition of neuromorphic models that are best captured as dynamical systems in continuous time. Any representation of a continuous system as a sequence of digital instructions implies fusing it with numerical integration techniques, thus intertwining the integration with the desired dynamics. Additionally, neuromorphic simulators and hardware platforms operate on the level of neuronal dynamics, a fundamentally different abstraction than digital intermediate representations for conventional machine instruction set architectures. This renders existing intermediate representations such as ONNX and MLIR poorly suited to describe neuromorphic computations faithfully.
In a joint effort between academia and industry, we present the Neuromorphic Intermediate Representation (NIR): an abstraction layer that bridges neuromorphic software and hardware platforms. NIR currently links 7 neuromorphic simulators (Lava, [10], Nengo [11], Norse [12], Rockpool [13], Sinabs [14], snnTorch [15], and Spyx [16]) and 4 neuromorphic hardware platforms (Loihi 2 [17] via Lava, Speck via Sinabs, SpiNNaker2 [18], and Xylo [19] via Rockpool). NIR defines graphs of computational nodes, where each node represents a computational primitive defined as a hybrid continuous-time dynamical system. This idealized description provides three distinct advantages compared to digital representations: (1) it avoids any assumptions around discretization or hardware constraints, (2) it provides a normative description against which implementations can be compared, and (3) it paves the way for seamless integration with mixed-signal chips, as well as analog accelerators and substrates. The computational graph representation is shown in Figure 1 along with its relation to digital and analog hardware as well as software, independently of the underlying continuous-time or discrete-time execution model.
In this paper, we (1) derive and implement a set of computational primitives that are common to the varied neuromorphic software and hardware landscape, (2) provide support for NIR across 7 different simulators and 4 different hardware platforms, (3) demonstrate flexible cross-platform deployment of NIR models, and (4) evaluate exemplary neuromorphic models defined in NIR across all 11 supported platforms—hardware and software. We develop our models on software simulators that are external and unfamiliar to the various neuromorphic hardware platforms we later execute the models on. The deployment of software-generated models on neuromorphic hardware was previously only possible through hardware-specific software libraries, and NIR effectively decouples the development of neural models from platform-specific tools. NIR now allows users to leverage tools and resources from the active neuromorphic open-source community and allows model development and platform-specific tools to develop efficiently in parallel, as illustrated in Figure 2). In this way, NIR increases accessibility to—and flexibility of—neuromorphic hardware platforms. This work comes at a decisive moment for novel compute technologies, where the power consumption of neural networks grows at an unprecedented rate, but also at a time when the platforms become available on a consumer scale for the first time.
1.1 Related work
Several works have called for widening usability and access to neuromorphic hardware and simulators [4, 20, 21] through co-design [5, 22] and shared representations [23]. PyNN [24] allows users to define neuroscientific spiking neural network (SNN) models in a simulator-independent language. PyNN models can be run without modifications on different neural simulators (NEST, Neuron, Arbor, Brian2) or neuromorphic hardware (e.g., BrainScaleS or SpiNNaker). PyNN is a programmatic frontend and does not provide a serializable data format. This is addressed by NeuroML [25] which supplies standards for defining and exchanging biological cell and network models. Both PyNN and NeuroML focus on computational neuroscience and not on the scalable compute capabilities of SNNs.
Several frameworks have been developed to design neuromorphic algorithms. Fugu [26] provides a high-level API to define and combine spiking neural algorithms. It generates an SNN graph as an intermediate representation which can serve as input to neuromorphic simulators or hardware compilers. Similarly, Lava [10] is a framework for developing brain-inspired neural network models and mapping them to Intel’s digital neuromorphic hardware. Its intermediate representation implements the Communicating Sequential Processes (CSP) principle [27], where neurons are represented by processes which send spikes via channels connected to other processes. Lava significantly simplifies the programming of the neuromorphic chip Loihi 2 [28]. However, at its core, it requires synchronization messages for triggering neuron updates, which is incompatible with analog neuromorphic systems. In addition to Fugu and Lava, there are frameworks such as the SNN-Toolbox [29] or NxTF [30] for converting neural network models from ML frameworks to specific SNN simulators and neuromorphic hardware. Zhang et al. [31] have developed an abstraction hierachy for brain-inspired computing with loose guarantees, portable across both von Neumann and neuromorphic architectures. Other interfaces focus on particular computing elements, such as crossbar arrays using beyond-CMOS technologies [32, 23, 33], or ReLU and leaky integrate-and-fire units [34]. Unfortunately, none of these standards have spread beyond their intended subdomain within neuromorphic computing.
Figure 2: a-d) shows four NIR primitives, where the name of the primitive is highlighted on top, the implemented computation is illustrated in the white box, and the parameters that are stored with the NIR graph are highlighted on the bottom. a) shows a stateful primitive, while b) and c) show stateless primitives, and d) shows a higher-order primitive. e-f) illustrate the concept of an intermediate representation (IR) with 7 software (SW) and 4 hardware (HW) backends. Instead 30 different compilers covering all $m\times n$ cases (e)), only $m+n$ interfaces between each hardware and software platform to NIR are necessary (f)).
2 Results
The Neuromorphic Intermediate Representation (NIR) uses signal-flow graphs to describe the intricate dynamics of neuromorphic neural networks. NIR captures coupled hybrid systems in a structured graph representation, shown in Figure 1 and 2, which are formally introduced in Section 2.2. In contrast to existing representations of neural dynamics, NIR operates on a predefined set of computational primitives rather than custom neural equations, which cannot easily be supported on efficient and specialized hardware. This approach is motivated by the nearly universal agreement on a few neuron models, such as the leaky integrate-and-fire model, in prevalent neuromorphic hardware platforms and simulators. The reductionist approach allows us to achieve strong generality early on; NIR is presently supported by four digital neuromorphic hardware platforms: Intel Loihi 2 [10], SynSense Speck, SpiNNaker2 [18], and SynSense Xylo [19]. Additionally, 7 software simulators now support NIR: Lava and Lava-DL, [10], Nengo [11], Norse [12], Rockpool [13], Sinabs [14], snnTorch [15], and Spyx [16].
The primary strength of NIR resides in its ability to reproduce the same computation on heterogeneous platforms, abstracting away platform-specific intricacies. Below, we demonstrate how a NIR graph can be executed and supported by the above-mentioned platforms across three heterogeneous tasks: a leaky integrate-and-fire model, a spiking convolutional network, and a recurrent reservoir network. Later, in Sections 2.2 and 2.3, we establish the axioms underpinning NIR graphs and list the computational primitives we presently support.
It is important to note that the hardware constraints and implementation choices of each platform necessarily lead to divergent behaviors. NIR does not attempt to match the computation perfectly on every platform. Rather, NIR aims to establish an idealized model that serves as a common ground for comparative analysis while highlighting the inter-platform discrepancies. By disentangling model dynamics from hardware idiosyncrasies, NIR encourages the development of platform-independent models that can subsequently be tuned to a specific platform, using, for instance, quantization-aware training. Our contribution serves a broader objective of neuromorphic computing: to simplify the creation of biologically inspired efficient neural networks, facilitating their consistent functionality across a wide range of computational environments.
2.1 Experiments
We conduct a series of experiments spanning diverse neuromorphic platforms to gauge the robustness and versatility of NIR. We chose three distinct tasks that are common within neuromorphic computing, through which we assess and compare performance across all compatible platforms: a leaky integrate-and fire neuron, a spiking convolutional neural network, and a recurrent reservoir model (see Figure 3).
Figure 3: The experiments we evaluate across 11 neuromorphic platforms and example data from the NMNIST and Braille dataset. (a) A single leaky integrate-and-fire neuron model, (b) a recurrent Braille classification model using current-based leaky integrate-and-fire in a spiking recurrent neural network (SRNN), and (c) a spiking convolutional neural network (SCNN). (d) and (e) shows sample data for the NMNIST and Braille datasets, respectively. The NMNIST activity is averaged across 300 timesteps.
A central difficulty in transferring a computation from one neuromorphic platform to another is that the neuron models and discretization schemes do not always match up exactly. There is a large amount of variation between differently implemented neuron models, even for the seemingly simple leaky integrate-and-fire (LIF) neuron. A more detailed discussion of the subtle differences in neuron models supported by NIR is presented in Section 4.4.
2.1.1 Leaky integrate-and-fire dynamics
As a first experiment, we demonstrate the behavior of a single leaky integrate-and-fire neuron under an identical input spike train (Figure 4 (a)) for every available platform. The LIF neuron was first instantiated on the software platform Norse, then exported into a NIR graph, and subsequently imported and executed by all different platforms. Figure 4 (a) visualizes both the voltage traces and output spikes from the platforms.
Figure 4: Experimental results. (a) Single leaky integrate-and fire neuron ordered from top to bottom: input spikes, voltage traces, and output activations. The timestep of spikes are well-aligned and only differ systematically for Rockpool and Sinabs due to discretization differences. Membrane potentials are normalized to lie between the resting potential $v_{rest}=0$ and the firing threshold $\vartheta$ to ignore platform-specific details regarding the numerical representations. (b) Spiking convolutional neural network: a platform-by-platform comparison of the spiking activity from the first spiking layer using cosine similarity (1 equals perfect overlap). Sinabs, snnTorch and Speck deviate from most other implementations due to their discretization choices. (c) and (d) Spiking recurrent neural network: similarity measure between the spiking activity of the first CuBa-LIF layer for an SRNN with biases and reset to zero and an SRNN without biases and subtractive reset.
${}^{\dagger}$ Nengo Norse Rockpool Sinabs snnTorch Speck SpiNNaker2 Spyx Xylo
SCNN 98.2% 98.1% 98.1% N/A 98.5% 97.9% 95.4% 98.2% 97.1% N/A
SRNN (Subtr.) N/A 78.6% 93.57% 71.4% N/A 92.1% N/A 93.57% 92.12% 85.71%
SRNN (Zero) 48.6% 55.7% 94.29% N/R N/A 95.0% N/A 85.00% 84.29% N/A
Table 1: Experimental accuracies for the spiking convolutional neural network (SCNN) and the spiking recurrent neural network (SRNN) measured against unseen test dataset. ${}^{\dagger}$ Lava offers a bitwise accurate simulator for Loihi 2, which is why the Loihi 2 column is merged with Lava.
To assess the viability of NIR as an intermediate representation for neuromorphic computing, we ensure that the dynamics are consistent across all platforms regarding the most common ways of encoding information in spiking neural networks—namely, rate encoding and spike time encoding [15]. The equivalence of output firing rates across different platforms can easily be seen, as the number of output spikes is identical across all platforms and since the spiking timings mostly coincide. Rockpool and Sinabs notably differ because of their integration scheme. They integrate the incoming signal ($v^{-}$, equation 1a) when checking for a threshold crossing before they update the leak term ($v^{+}$, equation 1b), whereas most other simulators immediately integrate the leak (equation 2). This effectively brings the neuron closer to the firing threshold earlier on, as observed in Figure 4 (a).
$v^{-}(t+1)$ $=\frac{dt}{\tau}\left(v(t)+v_{\text{leak}}+Ri(t+1)\right)$ (1a)
$v^{+}(t+1)$ $=v(t)-v^{-}(t+1)$ (1b)
$v(t+1)$ $=\frac{dt}{\tau}\left(v(t)+v_{\text{leak}}+Ri(t+1)\right)$ (2)
We further observe that the spike times are systematically delayed by one timestep for Lava and Loihi 2. This is because Lava calculates whether a neuron spikes at timestep $t$ based on its membrane potential at timestep $t-1$, i.e., before the addition of the input to the membrane, rather than after the input was added to the membrane potential. This systematic difference will only have an effect if absolute spike timing is used, but will have no impact if relative spike timing is used, that is if information is encoded in the timing between spikes.
Xylo and Speck are omitted, since they do not support the leaky integrate-and-fire primitive.
2.1.2 Convolutional neural network
For the second experiment, we studied a 9-layer spiking convolutional neural network (SCNN) using 2-dimensional convolutions, sum pooling, and integrate-and-fire neurons, shown in Figure 3 (c). The accuracies of the SCNNs across the different platforms on the withheld test dataset are shown to be very similar, see Table 1, with a mean test accuracy of 98.1% and a standard deviation of only 0.1% across all platforms. These results demonstrate that NIR can effectively act as an intermediate representation of a trained, rate-based spiking convolutional neural network, as the end-task performance is preserved across all tested platforms.
We further analyze the activities of the SCNN’s hidden layers across all different platforms. For a visually interpretable comparison, we compare the time-averaged activities for the first channel of the first hidden layer of I&F neurons in Figure 4 (b). The similarity measure tells us that most platforms have almost identical dynamics, while particularly Rockpool, Sinabs, snnTorch, and Speck deviates. Since the accuracies are relatively close, we observe that the rate encoding and feed-forward architecture of the investigated SCNN are relatively robust to these mismatches.
Rockpool and Xylo are omitted from this comparison because neither support convolutions, and Xylo does not implement the IF neuron model used in the SCNN.
2.1.3 Recurrent neural network
For the third and final task, we studied a two-layer spiking recurrent neural network (SRNN), with populations of current-based LIF (CuBa-LIF) neurons. The network was trained in snnTorch with backpropagation-through-time (BPTT) and surrogate gradients to perform a Braille letter recognition task [35]. After training, the network was exported to NIR, and evaluated on all other platforms. The accuracies and similarity measures for their respective activations are shown in Figure 4. We train two different networks, depending on the discretization of the reset mechanisms for post-spike membrane reset (shown in Figure 2 and equations (5) and (6)). Note that the NIR graph structure is the same, only the parameter values differ.
In this experiment, the performance across different platforms was notably less consistent than in the prior two tasks. Contrarily, the SRNN presented here displayed significant disparities in end-task performance when implemented on different platforms. This divergence can be attributed to the recurrent connections within the network, which tend to accentuate subtle discrepancies in the dynamics. The SRNN’s training on snnTorch suggests that its dynamics are intricately calibrated to the specificities of this simulator. Therefore, deploying the network on a platform with dissimilarities from snnTorch’s nuances can introduce pronounced differences in the system’s dynamics, thereby affecting the end-task outcome.
Hence, NIR emerges not only as a representation but also as an instrumental tool—it facilitates a deeper exploration of the discrepancies between platforms, revealing the influence of diverse parameterization, levels of precision, and models on the resultant performance mismatches. NIR thus offers a valuable avenue for research into the robustness of neural networks, specifically gauging their resilience to simulator and device mismatch, and thereby stands as a pivotal tool for the advancement of neuromorphic computing.
2.2 NIR axioms
The nodes in NIR are described by hybrid continuous-time systems, capturing the evolution of some physical quantity as well as jumps and discrete resets. Rather than executing these dynamics, the NIR graph merely declares them, ensuring that any computational backend is equipped with complete information required to interpret and evaluate the represented computation. It is essential to clarify that NIR’s objective is not to rectify discrepancies arising from mismatches between platforms. Rather, NIR offers a common ground upon which all hardware platforms can address shared challenges and brings into sharp focus the inconsistencies that arise from a heterogeneous neuromorphic computing landscape where neuron models and discretization schemes are not standardized.
Axiom 1 (Neuromorphic computational primitives)
NIR provides a set of standardized computational primitives $\mathcal{C}$. Each computational primitive $c\in\mathcal{C}$ defines a parameterized transformation $T_{\theta}$ of a continuous-time input signal $u(t)$ into a continuous-time output signal $y(t)$, where $\theta$ is the set of parameters for a given compute primitive $c$.
  1. (1.1)
    Every computational primitive $c$ is modeled as a hybrid system that describes how the computational dynamics evolve over time. This includes differential equations ($\dot{x}=f_{\theta}(x,u)$) as well as discretized transition functions for sudden activations, where $x(t)\in\mathbb{R}^{N_{x}}$, $u(t)\in\mathbb{R}^{N_{u}}$ are multidimensional signals. For brevity, we represent the set of parameters for a particular compute primitive with $\theta$, although this set of parameters varies for every primitive (see Table 2). For example, a leaky integrator may be modeled with the ODE $\dot{x}=-x+u$, and a linear connection layer may be modeled with the transfer function $y=Wx$.
  2. (1.2)
    Every primitive offers a set of input ports $p_{\text{in}}\in\mathcal{P}$, and a set of output ports $p_{\text{out}}\in\mathcal{P}$. A single port consists of a name and the expected shape of signal, $\mathcal{P}=(\mathcal{S}\times\mathbb{N}^{n})$ where $\mathcal{S}$ is the set of strings and $n$ is the number of independent signal channels. By exposing one or more ports, a single computational primitive can receive multiple input signals $u(t)=(u_{1}(t),\ldots,u_{n_{u}}(t))$ and send multiple output signals $y(t)=(y_{1}(t),\ldots,y_{n_{y}}(t))$, effectively separating semantically different signals. For example, a neuron may receive one input signal that models a synaptic current input, and another input signal that serves as a neuromodulatory signal. Ports are uniquely identifiable via a name, which, in the case of input ports, is similar to the argument name of a function.
  3. (1.3)
    Computational primitives may be composed into higher-order primitives, as shown in the bottom of Table 2 and illustrated in Figure 2.
Axiom 2 (Graph-based computation)
Every computation is represented as a NIR graph $\mathcal{G}=(\mathcal{V}\times\mathcal{E})$ where $\mathcal{V}$ is the set of computational nodes (e.g., a leaky integrate-and-fire neuron, or a linear connection layer), and $\mathcal{E}$ is the set of directed connections between the nodes in the graph.
  1. (2.1)
    A computational node $v\in\mathcal{V}$ is composed of a primitive $c\in\mathcal{C}$ and its parameters $\theta$. This yields a complete description of an input-output transformation, and thus fully describes the computation implemented by the computational node $v$. For example, the computational node for an affine connection layer consists of the primitive $c_{\text{Affine}}$ along with the weight matrix $W$ and the bias $b$ as the parameters $\theta=(W,b)$, such that $v_{\text{Affine}}=(c_{\text{Affine}},\theta)$.
  2. (2.2)
    An edge is a directional connection from an outgoing port to an incoming port $\mathcal{E}=(\mathcal{V}\times\mathcal{P}_{\text{in}})\times(\mathcal{V}\times\mathcal{P}_{\text{out}})$, where $\mathcal{P}_{\text{in}}$ and $\mathcal{P}_{\text{out}}$ are the ports of the input and output nodes, respectively. There are no restrictions on the number of connections any node or port can have. As an example, the edge $e=((v_{\text{Affine}},p_{\text{out}}),(v_{\text{Affine}},p_{\text{in}}))$ connects the output of the affine node to the input of that same node.
  1. (2.2.1)
    Edges do not perform any computation, and are effectively the identity map $I:x\mapsto x$.
  2. (2.2.2)
    One input port can receive multiple incoming edges which are then, by convention, summed together elementwise. Similarly, one output port can send its information via different edges to various input ports. The information is then passed by values, not by reference, such that information can only flow along the direction of an edge.
  1. (2.3)
    Every NIR graph has exactly one input node and one output node. These nodes do not transform their input—they exist to disambiguate where the input and output of a graph is situated.
The above axioms closely follow conventional descriptions of directed signal-flow graphs. Figure 2 further illustrates the workings of NIR graphs, and Figure 3 shows three example NIR graphs that are used in the experiments from Section 2.1.
2.3 Computational primitives
Primitive Parameters Parameter types Computation
Input
Input shape
$\mathbb{N}^{N_{0}\cdots N_{n}}$
Output
Output shape
$\mathbb{N}^{N_{0}\cdots N_{0}}$
Affine
$W,b$
$\mathbb{R}^{N_{0}\cdots N_{n}}\times\mathbb{R}^{N_{1}\cdots N_{n}}$
$W\,i(\cdot)+b$
Convolution
$W$, Stride, Padding, Dilation, Groups, Bias
$\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}\times N_{0}\cdots N_{n}}\times\mathbb{N}^{N_{0}\cdots N_{0}}\times\mathbb{N}^{N_{0}\cdots N_{0}}\times\mathbb{N}^{N_{0}\cdots N_{0}}\times\mathbb{N}\times\mathbb{R}^{N_{0}\cdots N_{0}}$
$f\star g$
Delay
$\tau$
$\mathbb{R}^{N}$
$i(t-\tau)$
Flatten
Input shape, Start dim., End dim.
$\mathbb{N}^{N_{0}\cdots N_{n}}\times\mathbb{N}\times\mathbb{N}$
Integrator
$R$
$\mathbb{R}^{N_{0}\cdots N_{n}}$
$\dot{v}=R\,i(\cdot)$
Leaky integrator (LI)
$\tau$, $R$, $v_{leak}$
$\mathbb{R}^{N_{0}\cdots N_{n}}\times\mathbb{R}^{N_{0}\cdots N_{n}}\times\mathbb{R}^{N_{0}\cdots N_{n}}$
$\tau\dot{v}=(v_{leak}-v)+R\,i(\cdot)$
Linear
$W$
$\mathbb{R}^{N_{0}\cdots N_{n}}$
$W\,i(\cdot)$
Scale
$s$
$\mathbb{R}^{N_{0}\cdots N_{n}}$
$s\,i(\cdot)$
Spike
$\theta_{\text{thr}}$
$\mathbb{R}^{N_{0}\cdots N_{n}}$
$\delta(i(\cdot)-\theta_{\text{thr}})$
Higher-order primitive Computation
[…]
Table 2: The computational primitives, their parameters and computational models (top). The bottom listing defines higher-order primitives as compositions. Parameters are typed for an arbitrary number of dimensions, $n$. $i(\cdot)$ denotes input at a given point in time, where $t$ denotes a specific time, $C_{\text{out}}\times C_{\text{in}}$ denotes convolutional output and input channels, and $\delta$ denotes the Dirac delta function.
NIR defines 11 computational primitives and 3 higher-order primitives, listed in Table 2. We define common neuromorphic components like the linear map, leaky integrator, and spike threshold function, but we further included mathematical primitives such as the affine map and convolution. The input and output nodes serve to disambiguate the entries and exits of a graph.
The 11 primitives in NIR are “fundamental” in the sense that the backends implementing the primitives are required to approximate the computation of the idealized description as closely as possible within the limitations of the platform. Any given platform is not expected to implement the full specification. This is particularly true for functionally specialized hardware, where hardware restrictions render certain functional primitives impossible. We detail such restrictions in Section 4.3.1.
NIR does not restrict the implementation apart from stating the idealized descriptions, and it is expected that the exact computation of the primitives will vary across platforms. This is particularly important for digital systems and their choice of integration and discretization method. Given the heterogeneity of the present hardware landscape, this amount of freedom becomes a practical choice as well; any restrictions towards discretizing the hybrid ODEs will invariably lead to incompatibilities and numerical deviations across platforms.
The primitives listed in Table 2 can be composed to provide more complex computational elements. Three such examples are already present in the specifications as convenient abstractions:
  1. 1.
    The integrate-and-fire (IF) neuron, defined as an integrator composed with a threshold and membrane reset mechanism.
  2. 2.
    The leaky integrate-and-fire (LIF) neuron, defined as a leaky integrator composed with a threshold and membrane reset mechanism.
  3. 3.
    The current-based leaky integrate-and-fire neuron (CuBa-LIF) defined as a leaky integrator, linearity, and leaky integrate-and-fire neuron.
3 Discussion
We presented an intermediate representation for continuous-time neuromorphic systems, dubbed the Neuromorphic Intermediate Representation (NIR), in which we established a set of primitives that can be arbitrarily composed to create sophisticated computational structures. We demonstrated the application of NIR by evaluating three tasks across the supported neuromorphic platforms: a leaky integrate-and-fire model, a spiking convolutional neural network using integrate-and-fire neurons, and a spiking recurrent neural network using current-based leaky integrate-and-fire neurons. We found that NIR faithfully represents the idealized underlying computation and that the execution aligns well across the supported platforms for feed-forward dynamics, such as a single LIF neuron or a spiking convolutional network. For more complex and time-sensitive models, such as the recurrent network, we found that platform-specific constraints and discretization choices produced somewhat diverging activation patterns and accuracies.
The possibility of conversion between different software systems allows a user to leverage features of different libraries such as quantization aware training or particularly fast training speeds. This cross-compatibility paves the way for a versatile, shared SNN tooling infrastructure.
Because NIR represents an idealized model, we argue that NIR helps highlight differences between neuromorphic hardware and software that would be difficult to identify otherwise. The common representation in NIR therefore not only permits the translation of neuromorphic computational models between systems, but also simplifies platform-specific optimizations. While such optimizations may reduce the ability for the model to generalize, it is an important component if neuromorphic systems are to compete with conventional compute platforms.
One crucial implication that we do not discuss in this work is the connection to physical and analog systems. The primitives in NIR are effectively first-order integrators and can be composed to express higher-order primitives, as in the LIF and CuBa-LIF models. In turn, this can be used to express more complex partial differential equations and model problems like diffusion, fluid dynamics, and quantum mechanics. Effectively, our method generalizes to other kinds of hardware, including analog substrates.
Our main goal has been to address the lack of shared representations for neuromorphic computing. NIR effectively guarantees reproducibility across neuromorphic simulators and hardware. Similar to how the initial design of digital instruction sets propelled the advent of modern-day computing, we believe this to be a crucial step towards further development of neuromorphic computing. We hope that NIR can accelerate both the academic development required to propel neuromorphic computing to perform on-par with present-day deep learning, while solving practical problems in the near future.
NIR is open-source, licensed under BSD-3, and available at https://github.com/neuromorphs/NIR
4 Methods
4.1 Computational graphs
Computational graphs are interconnected computational nodes through which signals flow, and have been an important tool for computational studies in analog and digital systems since their inception in the 1940s [36]. In its most abstract form, computational graphs ($\mathcal{G}\coloneqq(\mathcal{V},\mathcal{E})$) consist of a set of nodes ($\mathcal{V}$) and edges ($\mathcal{E}\coloneqq(\mathcal{V}\times\mathcal{V})$) connecting two nodes. An example of a computational graph could be discrete components on a circuit board, where wires connect components together. Any state needed in the individual components, such as voltage, memory, etc., would be dealt with internally. Another example could be layers in a machine learning model, where the order in which the layers are applied determines the flow of the input signal through the model. Most machine learning models do not require state contrary to the circuit example, although some layers do, such as batch normalization or recurrent units.
[…]
$T$$O$$X$$S$$Y$
Figure 5: Left: a signal flow graph where a signal $X$ is integrated by some arbitrary process to produce the output signal, $Y$. Right: a Mealy machine where a transition node $T$ updates a recurrent state $S$ that is forwarded to an output node $O$ who, in turn, determines the output $Y$.
In the 1950’s Mealy introduced a notation that explicitly captures state, as it would have to be stored in, say, digital circuits [37]. Figure 5 visualizes the Mealy machine on the right side, where some input signal, $X$, is transformed according to some ruleset in $T$ that recursively operates on the state $S$. Both $S$ and $X$ are then transformed according to another ruleset $O$ which computes the final output $Y$. Mealy machines formally separate memory (state) from computation (nodes). The same distinction cannot be made in analog systems, where the system itself represents the state and computation. While signal flow graphs are technically more ambiguous regarding the representation of state, they compactly represent computational graphs for both digital and analog systems.
In the following, we restrict ourselves to the domain of first-order systems, described by ordinary differential equations (ODEs) of the form
$\dot{x}=f(x;\theta),\quad x\in\mathbb{R}^{N},\ \ \theta\in\mathbb{R}^{M}$ (3)
where $x(t)$ is a real-valued function of time $t$ and $f(x)$ is a continuous real-valued function of the $N$-dimensional input $x$. $\theta$ describes a fixed set of $M$-dimensional parameters. $f$ may be smooth, but it may also be subject to sudden forces, such as a bouncing ball hitting a hard surface. For that reason, we use hybrid systems, defined as the combination of continuously smooth functions as well as functions with discrete transitions. We explicitly model jumps by allowing conditional differential equations of the form
$z=\begin{cases}f&\text{Condition}\\ g&\text{Else}\end{cases}$ (4)
where $f$ and $g$ are real-valued functions in time.
Digital systems require discrete instructions and cannot directly solve the equations above. Numerical methods for optimally discretizing continuous ODEs is a mature and active research area with a long history [38]. Despite the discrepancy between the continuous-time ODE formulation and digital computers, modern fixed-point neural solvers allow for sound accuracy with relatively large timesteps and low precision [39]. Therefore, we argue that the continuous-time ODE formulation is not a hindrance for digital systems. In fact, it allows NIR to express computations at a higher level of abstraction that is independent of how the system is discretized. This results in inaccuracies across systems that we address in Section 2.3.
4.2 Modeling hardware using NIR
NIR provides a framework to represent neuromorphic hardware as specific subclasses of NIR graphs, i.e., the subset of all NIR graphs that are compatible with the given hardware system. This representation allows direct translation between any simulation-defined computation and a target hardware platform.
Each hardware system imposes inherent constraints on executable computations. For instance, the Xylo Audio 2 chip limits the network to a maximum of 1000 LIF neurons, with each neuron supporting two synaptic states and a maximum fan-out of 64. Moreover, networks on Xylo can have only one hidden layer of LIF neurons, with input weights, recurrent weights, and output weights to an output LIF neuron population of size $\leq 8$. Such constraints can be codified into a defined set of NIR graphs, denoted as $\mathbb{G}_{xylo}\subset\mathbb{G}$. For any prospective NIR graph $g\in\mathbb{G}$, determining its compatibility with the Xylo Audio 2 is a matter of verifying $g$ against $\mathbb{G}_{xylo}$.
4.3 Neuromorphic simulators and hardware platforms
NIR integrates with a number of simulators and hardware platforms that implement the de facto dynamics of the underlying hybrid ODEs. Specifically, they have made a series of choices around the numerical integration of the systems of equations that we detail alphabetically below.
4.3.1 Hardware platforms
Loihi 2
The Loihi 2 chip by Intel consists of 6 embedded microprocessor cores (Lakemont x86) and 128 fully asynchronous neuron cores (NCs) connected by a network-on-chip. The NCs are optimized for neuromorphic workloads by implementing a group of spiking neurons and including all synapses connected to such neurons. All the communication between NCs is in the form of spike messages. Microprocessor cores are optimized for spike-based communication and execute standard C code to assist with data I/O as well as network configuration, management and monitoring. Some of the new functionalities added in this second version of the Loihi chip are the possibility of implementing custom neuron models using microcode instructions (assembly), the option to generate and transmit graded spikes, and support for three-factor learning rules. A single Loihi 2 chip supports up to 1 million neurons and 120 million synapses. Together with Loihi 2, Intel presented their open-source framework Lava, that allows users to write neuro-inspired applications and map them to both traditional and neuromorphic hardware. Using high-level Python APIs, users can describe their neural networks, which are then compiled to run on the requested backend. Currently, Lava supports deployment on traditional CPU and Loihi 2. Specifically for Loihi 2, Lava also gives the possibility of writing custom neuron models in assembly to be run on the NCs, and custom C code to be run in the LMTs.
SpiNNaker2
SpiNNaker2 is an IC designed for simulation of very large scale spiking neural networks. In contrast to many other solutions, it uses 152 processing elements (PE) connected by a network on chip. Each PE contains an ARM Cortex M4F core with 128 kB local SRAM, as well as accelerators for exponential and logarithm functions and a 16x4 MAC array for 2D convolution and matrix-matrix multiplications [40]. Multiple chips can be connected using dedicated chip2chip links and an on-chip packet router optimized for small packets of spikes. In addition, each chip can be connected to LPDDR4 DRAM in order to extend the amount of memory available per chip. The py-spinnaker2 software framework [41] uses 8-bit signed synapse weights and 32-bit floating-point numbers for neuron parameters and state variables. Currently, IAF, LIF and CuBa-LIF neuron models are implemented and support both reset by subtraction or reset to zero. Since SpiNNaker2 is software-based, further models can be added. It is also not limited to SNN execution, but can be used for any computational task, including real-time control or deep neural networks.
SynSense Speck
Speck is an integrated sensor-processor IC that fuses event-based vision sensing with event-driven spiking CNN processing. The ultra-low-power IC operates fully asynchronously and takes full advantage of the asynchronous nature of events produced by the DVS. These events are processed by Integrate and Fire neurons that are interconnected efficiently using a convolutional engine within each core. In the chip version used in this work comprises 9 dedicated SCNN cores, each consisting of convolutional connections, IAF neurons and pooling. The chip supports 8 bit weight resolution and 16 bit membrane resolution. Each core supports convolutions of up to 1024 input and output channels with strides of 1, 2, 4 or 8. The device is accessible via a high-level Python library Sinabs or a low-level library samna. Further details are available at sinabs.ai and synsense.ai.
SynSense Xylo
Xylo is a series of ultra-low-power devices for sensory inference, featuring a digital SNN core adaptable to various sensory inputs like audio and bio-signals. Its SNN core uses an integer-logic CuBa-LIF neuron model with customizable parameters for each synapse and neuron, supporting a wide range of network architectures. The Xylo Audio 2 model (SYNS61201) specifically includes 8-bit synaptic weights, 16-bit synaptic and membrane states, two synaptic states per neuron, 16 input channels, 1000 hidden neurons, 8 output neurons with 8 output channels, a maximum fan-in of 63, and a total of 64,000 synaptic weights. For more detailed technical information, see https://rockpool.ai/devices/xylo-overview.html. The Rockpool toolchain contains quantizaton methods designed for Xylo, as well as bit-accurate simulations of Xylo devices.
4.3.2 Simulation frameworks
Most simulation frameworks mentioned here are based on the machine learning accelerator PyTorch [42]. PyTorch effectively operates as a compiler that translates Python code to various digital compute architectures, including CPUs and GPUs. This does not absolve the simulation frameworks from addressing the discretization problems mentioned above, although they can freely use PyTorch-related features such as quantization and varying floating-point precision to approximate hardware constraints.
Lava
is an open-source software stack developed by Intel and used for programming their Loihi 2 chip [10]. Lava has a modular structure, supporting versatile processes from neurons to external device interfaces. These processes communicate via event-based messaging and are adaptable to various platforms, including CPU, GPU and the Loihi 2 chip. Within Lava, there are several other sub-libraries. For NIR, we are using Lava itself as well as Lava-dl, which offers offline training, online training and inference methods for various deep event-based networks.
Nengo
is used to implement networks for deep learning, vision, motor control, visual attention, serial recall, action selection, working memory, attractor dynamics, inductive reasoning, path integration, and planning with problem-solving [11]. Nengo has libraries specifically designed to help with cognitive modelling, deep learning, adaptive control, and accurate dynamics, among others. While Nengo aims to be agnostic to particular neuron models by automatically locally retraining weights, here we use the neuron model of its default CPU implementation.
Norse
models stateless networks by explicitly passing the state of the neuron computation into each computation [12]. This approach simplifies parallelization and sharding for both the model and the data. Norse applies simple feed-forward Euler integration and implements various surrogate and adjoint methods for gradient-based optimization, as well as spike-time dependent plasticity and Tsodyks-Markram short-term plasticity for unsupervised adaptation.
Rockpool
is a machine-learning framework for SNNs, supporting network design, training, testing, and deployment to neuromorphic hardware [13]. Rockpool provides a torch-like interface, with automatic differentiation back-ends and hardware acceleration based on PyTorch and JAX[43]. The library includes hardware-aware training, including quantization- and pruning-in-training, as well as post-training quantization. Rockpool includes a flexible and extensible deployment framework based on graph extraction, which currently includes deployment to multiple devices in the Xylo family.
Sinabs
is a deep learning library based on PyTorch for spiking neural networks, with a focus on simplicity, fast training and extensibility [14]. Sinabs works well for computer vision models as it supports weight transfer from conventional CNNs and enables deployment to Speck, the spiking convolutional processor. With its support for EXODUS [44], it allows for extremely fast training of deep SNNs. It integrates seamlessly into libraries built on top of PyTorch such as Lightning.
snnTorch
provides a thin abstraction layer on top of PyTorch for training and modelling SNNs [15]. It prioritizes flexibility by providing the option of stateless and stateful networks to the user. It integrates various learning rules including customizable surrogate gradient descent with backpropagation through time, along with online learning rules such as real-time recurrent learning variants and spike-time dependent plasticity. snnTorch also accounts for hardware-friendly training approaches including stateful quantization that digital accelerators may need to account for, as well as probabilistic neuron models that factor in noise models typical of analog or mixed-signal hardware.
Spyx
[16] is a light-weight and modular package for training SNNs within the JAX ecosystem [43]. Extending Google Deepmind’s Haiku library [45] for training deep neural networks, Spyx offers a host of simplified spiking neuron models and utility functions that make it easy to compose and train SNN architectures through surrogate gradient descent or neuroevolution. A notable feature is the ability for users to utilize mixed-precision with minimal code modification and leverage Just-In-Time (JIT) compilation for the entire training loop to achieve maximal hardware utilization on modern deep learning accelerators such as GPUs or TPUs.
4.4 Neuron model parameterizations
We now turn to the implementation details of NIR equations across the different platforms, with a particular focus on leaky integrate-and-fire dynamics from experiments 1 and 3 in Section 2.1. We note that the idealized reset can be discretized by either a hard reset to $\theta_{\text{thr}}$ or a “soft” subtractive reset that retains the residual from a possible overshoot:
Hard reset $v_{t+1}$ =$=\begin{cases}v_{t}&v_{t}<\theta_{\text{thr}}\\ 0&v_{t}\geqslant\theta_{\text{thr}}\end{cases}$ (5)
Subtractive reset $v_{t+1}$ =$=\begin{cases}v_{t}&v_{t}<\theta_{\text{thr}}\\ v_{t}-\theta_{\text{reset}}&v_{t}\geqslant\theta_{\text{thr}}\end{cases}$ (6)
The differences in the discretization of spiking neuron dynamics are detailed below; equations for Spyx are not listed, as they follow the same implementation as snnTorch.
4.4.1 LIF
NIR
The leaky integrate-and-fire model in NIR is given by:
$\tau\dot{v}=v_{\text{leak}}-v+Ri$ (7)
where $\tau$ is the time constant, $v_{\text{leak}}$ is the membrane leak, $R$ is the resistance and $i(t)$ is the input current. Discretized using Euler’s method, this turns into:
$v(t+1)=\left(1-\frac{dt}{\tau}\right)v(t)+\frac{dt}{\tau}v_{\text{leak}}+\frac{dt}{\tau}Ri(t+1)$ (8)
where $dt$ is the discretization timestep.
Exact solution for constant input current
The course of the membrane potential $v$ can be exactly calculated for a constant input current $i=i_{0}=\text{const.}$. and initial condition $v(0)=v_{0}$ [46, eq. 1.7]:
$v(t)=v_{\text{leak}}+Ri_{0}\left(1-\exp{\left(-\frac{t}{\tau}\right)}\right)+v_{0}\exp{\left(-\frac{t}{\tau}\right)}$ (9)
Lava-dl
In lava-dl, we used the CuBa-LIF model as:
$u(t)$ $=(1-\alpha_{u})u(t-1)+i(t)$ (10)
$v(t)$ $=(1-\alpha_{v})v(t-1)+u(t)$ (11)
with $\alpha_{u}=1$ to effectively turn it into a LIF model:
$v(t)$ $=(1-\alpha_{v})v(t-1)+i(t)$ (12)
with
$\alpha_{v}$ $=\frac{dt}{\tau_{\text{mem}}}$ (13)
With the constraints
$R\frac{dt}{\tau_{\text{mem}}}$ $=1$ (14)
$v_{\text{leak}}\frac{dt}{\tau_{\text{mem}}}$ $=0$ (15)
One may implement the first constraint ($R$) by hard-coding it into a linear layer, or into the voltage threshold. One may implement $v_{\text{leak}}$ through a bias term, but this was not done in our experiments.
Norse
The LIF model in Norse lacks the resistivity term, which we compensate for by scaling the input
$v(t+1)=v(t)+\frac{dt}{\tau}\left(v_{leak}-v(t)+i\right)$ (16)
snnTorch
The LIF model in snnTorch is given by:
$v(t+1)=\beta v(t)+i(t+1)$ (17)
where $\beta$ is the membrane potential decay rate. We have:
$\beta=1-\frac{dt}{\tau}$ (18)
And the constraints are:
$R\frac{dt}{\tau}$ $=1$ (19)
$v_{\text{leak}}\frac{dt}{\tau}$ $=0$ (20)
However, we can use a linear layer before the LIF in simulation to effectively implement the resistance term. The weight matrix of this linear layer is given by:
$W_{ij}=\begin{cases}R\frac{dt}{\tau}&\text{if }i=j\\ 0&\text{otherwise}\end{cases}$ (21)
SpiNNaker2
In py-spinnaker2 [41] the LIF model is given by:
$v(t)=\alpha_{\text{decay}}v(t-1)+i(t)+i_{\text{offset}}$ (22)
Two modes for the translation between NIR and SpiNNaker2 are supported:
In Exponential-Euler mode, the decay factor $\alpha_{\text{decay}}$ is calculated such that in case without input ($i=0$ and $i_{\text{offset}}=0$) and with $v_{\mathrm{leak}}=0$ the voltage exactly matches the continuous-time dynamics:
$v(t+dt)=\alpha_{\text{decay}}v(t)$ =!${=}v(t)\exp{(-\frac{dt}{\tau})}$ (23)
$\Rightarrow\alpha_{\text{decay}}$ $=\exp{(-\frac{dt}{\tau})}$ (24)
The SpiNNaker2 threshold $\Theta$ is scaled to cope for the non-exisiting resistance variable:
$\Theta=\frac{\theta_{\text{thr}}}{\left(1-\exp(-\frac{dt}{\tau})\right)R}$ (25)
In Forward-Euler mode we match the discrete Euler update of the LIF model (Eq. 8) so that decay factor and threshold become:
$\alpha_{\text{decay}}$ $=1-\frac{dt}{\tau}$ (26)
$\Theta$ $=\theta_{\text{thr}}\frac{\tau}{dtR}$ (27)
In both modes, the offset current $i_{\text{offset}}$ is used to consider the parameter $v_{\text{leak}}$ and the bias $i_{\text{bias}}$ from incoming Affine or Conv2D nodes:
$i_{\text{offset}}=\frac{v_{\text{leak}}}{R}+i_{\text{bias}}$ (28)
Nengo
The default LIF model in Nengo’s CPU implementation is identical to the Exponential-Euler model for SpiNNaker2 described above, except that it has a fixed threshold $\Theta$. To compensate for this we adjust the gain on each neuron (a multiplicative factor applied to all inputs).
Rockpool
The LIF model in Rockpool is a time-discretised neuron membrane, with update equations
$V_{mem}[n]=\alpha V_{mem}[n-1]$
$V_{mem}[n]=V_{mem}[n]+S_{in}[n]+b+\sigma\zeta[n]$
$V_{mem}[n]>\theta_{thr}\rightarrow S_{out}[n]=1$
$V_{mem}[n]=V_{mem}[n]-\theta$
where $\alpha=\exp(-dt/\tau_{mem})$. Here $V_{mem}$ is the floating-point membrane potential state; $S_{in}$ and $S_{out}$ are the incident and output spike trains respectively; $\sigma\zeta[n]$ is a scaled white noise process with mean 00 and std. dev. $\sigma$; $b$ is a bias current and $\theta$ is the firing threshold. We have neglected neuron indices in the above equations; all parameters are independently defined per neuron in Rockpool.
4.4.2 CuBa-LIF
NIR
The current-based LIF model in NIR is given by:
$\tau_{\text{syn}}\dot{u}$ $=-u+w_{\text{in}}i$ (29)
$\tau_{\text{mem}}\dot{v}$ $=v_{\text{leak}}-v+Ru$ (30)
where $u(t)$ is the synaptic current, $\tau_{\text{mem}}$ the membrane time constant, $\tau_{\text{syn}}$ the synaptic time constant, and $w_{\text{in}}$ the scaling factor of the current input $i(t)$. Discretized:
$u(t+1)$ $=\left(1-\frac{dt}{\tau_{\text{syn}}}\right)u(t)+w_{\text{in}}\frac{dt}{\tau_{\text{syn}}}i(t+1)$ (31)
$v(t+1)$ $=\left(1-\frac{dt}{\tau_{\text{mem}}}\right)v(t)+\frac{dt}{\tau_{\text{mem}}}v_{\text{leak}}+R\frac{dt}{\tau_{\text{mem}}}u(t)$ (32)
Lava-dl
In lava-dl, the model is:
$u(t)$ $=(1-\alpha_{u})u(t-1)+i(t)$ (33)
$v(t)$ $=(1-\alpha_{v})v(t-1)+u(t)$ (34)
Such that:
$\alpha_{u}$ $=\frac{dt}{\tau_{\text{syn}}}$ (35)
$\alpha_{v}$ $=\frac{dt}{\tau_{\text{mem}}}$ (36)
With the constraints
$w_{\text{in}}\frac{dt}{\tau_{\text{syn}}}$ $=1$ (37)
$R\frac{dt}{\tau_{\text{mem}}}$ $=1$ (38)
$v_{\text{leak}}\frac{dt}{\tau_{\text{mem}}}$ $=0$ (39)
Where the first constraint ($w_{\text{in}}$) can be hard-coded into a linear layer, or into the voltage threshold.
One may also implement $v_{\text{leak}}$ through a bias term, but this was not done in our experiments.
Norse
Norse discretizes the CuBa-LIF model as follows:
$u(t+1)$ $=u(t)-\frac{dt}{\tau_{syn}}i(t)$ (40)
$v(t+1)$ $=v(t)+\frac{dt}{\tau_{mem}}\left(v_{leak}-v+u\right)$ (41)
This deviates from the canonical formulation by lacking a weight and resistance term. The weight term is introduced by scaling the input, while the resistance term can be introduced by adding a weight factor between the synaptic ($u$) and somatic ($v$) parts.
snnTorch
$u(t+1)$ $=\alpha u(t)+i(t+1)$ (42)
$v(t+1)$ $=\beta v(t)+u(t+1)$ (43)
Leading to:
$\alpha$ $=1-\frac{dt}{\tau_{\text{syn}}}$ (44)
$\beta$ $=1-\frac{dt}{\tau_{\text{mem}}}$ (45)
With constraints:
$w_{\text{in}}\frac{dt}{\tau_{\text{syn}}}$ $=1$ (46)
$R\frac{dt}{\tau_{\text{mem}}}$ $=1$ (47)
$v_{\text{leak}}\frac{dt}{\tau_{\text{mem}}}$ $=0$ (48)
Where a linear layer can be used with:
$W_{ij}=\begin{cases}w_{\text{in}}\frac{dt}{\tau_{\text{syn}}}&\text{if }i=j\\ 0&\text{otherwise}\end{cases}$ (49)
to replace the first constraint.
snnTorch also offers a tunable membrane potential threshold which is shared across a population of neurons. As such, if $W$ is a multiple of the identity matrix, the first constraint can also be satisfied by simply scaling the membrane potential threshold accordingly.
SpiNNaker2
In py-spinnaker2 [41] the Cuba-LIF model is given by:
$I_{\text{syn}}(t)$ $=\alpha_{syn}I_{\text{syn}}(t-1)+i(t)$ (50)
$v(t)$ $=\alpha_{\text{mem}}v(t-1)+I_{\text{syn}}(t)+i_{\text{offset}}$ (51)
As for LIF, two modes for the translation between NIR and SpiNNaker2 are supported:
In Exponential-Euler mode, the decay factors, the threshold and $i_{\text{offset}}$ are calculated as follows:
$\alpha_{\text{mem}}$ $=\exp{(-\frac{dt}{\tau_{\text{mem}}})}$ (52)
$\alpha_{\text{syn}}$ $=\exp{(-\frac{dt}{\tau_{\text{syn}}})}$ (53)
$\Theta$ $=\theta_{\text{thr}}\left(\left(1-\exp(-\frac{dt}{\tau_{\text{mem}}})\right)R\left(1-\exp(-\frac{dt}{\tau_{syn}})\right)w_{\text{in}}\right)^{-1}$ (54)
$i_{\text{offset}}$ $=\frac{v_{\text{leak}}}{R}\left(\left(1-\exp(-\frac{dt}{\tau_{syn}})\right)w_{\text{in}}\right)^{-1}+i_{\text{bias}}\left(1-\exp(-\frac{dt}{\tau_{syn}})\right)^{-1}$ (55)
In Forward-Euler mode we match the discrete Cuba-LIF equations (40-41) so that SpiNNaker2 parameters become:
$\alpha_{\text{mem}}$ $=1-\frac{dt}{\tau_{mem}}$ (56)
$\alpha_{\text{syn}}$ $=1-\frac{dt}{\tau_{syn}}$ (57)
$\Theta$ $=\theta_{\text{thr}}\frac{\tau_{mem}}{dtR}\frac{\tau_{syn}}{dtw_{in}}$ (58)
$i_{\text{offset}}$ $=\frac{v_{\text{leak}}}{R}\frac{\tau_{syn}}{dtw_{in}}+i_{\text{bias}}\frac{dt}{\tau_{syn}}$ (59)
We note that in both modes the integration of the bias into $i_{\text{offset}}$ is not exact. Instead, the conversion is chosen such that for bias-only input the membrane voltage converges to the same value as in the NIR model.
Rockpool
The Rockpool CuBa-LIF neuron is a time-discretized simulation of exponential LIF dynamics.
$I_{syn}[n]=I_{syn}[n-1]+S_{in}[n]+W_{rec}\mathbf{S_{out}}[n-1]$
$I_{syn}[n]=\beta I_{syn}[n]$
$V_{mem}[n]=\alpha V_{mem}[n-1]$
$V_{mem}[n]=V_{mem}[n]+I_{syn}[n]+b_{j}+\sigma\zeta[n]$
$V_{mem}[n]>V_{thr}\rightarrow S_{out}[n]=1$
$V_{mem}[n]=V_{mem}[n]-V_{thr}$
where $\alpha=$ and $\beta=\exp(-dt/\tau_{syn})$. Here $W_{rec}$ is the recurrent weight matrix for the layer of neurons; $\mathbf{S_{out}}[n]$ is the vector of spikes produced by the layer at timestep $n$; and other symbols are as in the Rockpool LIF equations above. We have neglected neuron indices for simplicity; all neurons have independent state variables and parameters.
Xylo
The Xylo CuBa-LIF neuron is a time-discretized low-bit-depth integer logic neuron, simulating an approximation of exponential LIF dynamics using integer bit-shift decay circuitry. In the following, the state variables $I_{syn}$ and $V_{mem}$ are 16-bit signed integers; $W_{rec}$ are 8-bit signed integers; $d_{syn}$ and $d_{mem}$ are unsigned integer bit-shift decay parameters approximating $d\approx\log_{2}(\tau/dt)$; $V_{thr}$ and $b$ are 16-bit signed integers; $S_{in}$ and $S_{out}$ are unsigned integer number of events per $dt$.
$I_{syn}[n]=I_{syn}[n-1]+S_{in}[n]+W_{rec}\mathbf{S_{out}}[n-1]$
$I_{syn}[n]=\begin{cases}I_{syn}[n]-I_{syn}[n]>>d_{syn}&\text{if }I_{syn}[n]>>d_{syn}>0\\ I_{syn}[n]-1&\text{otherwise}\end{cases}$ (60)
$V_{mem}[n]=\begin{cases}V_{mem}[n]-V_{mem}[n]>>d_{mem}&\text{if }V_{mem}[n]>>d_{mem}>0\\ V_{mem}[n]-1&\text{otherwise}\end{cases}$ (61)
$V_{mem}[n]=V_{mem}[n]+I_{syn}[n]+b+\sigma\zeta[n]$
$V_{mem}[n]>\theta_{thr}\rightarrow S_{out}[n]=1$
$V_{mem}[n]=V_{mem}[n]-t\theta_{thr}$
where $x>>n$ is the right-shift operator, shifting $x$ by $n$ bits.
The bit-shift decay operation in Eqs. 60 and 61 approximates exponential decay for $V_{mem}$ and $I_{syn}$ state variables. When the integer quantised decay parameters $d_{*}$ obtained from $\tau_{*}$ are close to the corresponding floating-point values for $d_{*}$, this approximation causes only small deviations in dynamics. For values of $V_{mem}$ and $I_{syn}$ close to zero, where the bit-shift decay operation would not result in a decrease of the state variables, we include a unit decay per $dt$. We have neglected neuron indices in the above equations; all neurons have independent state variables and parameters.
4.5 Training setup
We proceed to describe the training setup used to obtain the architecture and parameters for the SCNN and SRNN in the second and third experiment, respectively. The training code can be found on https://github.com/neuromorphs/NIR/tree/paper/paper.
4.5.1 Spiking convolutional neural network
We trained a non-spiking ReLU based convolutional neural network on the neuromorphic MNIST dataset (N-MNIST) [47], that was later transferred to an equivalent spiking convolutional network with Sinabs, using the model transfer approach from  [48]. The last layer of the network contains 10 neurons, each one representing one digit, trained to have the largest firing rate for the neuron corresponding to the correct label. The training was done using backpropagation through time and surrogate gradients. Further details are available in our code repository.
4.5.2 Spiking recurrent neural network
Data from pressure readings encoded using a sigma-delta modulator with $\vartheta=1$ was used, accounting for a time binning with bin size equal to 5 ms. Compared to the original implementation, we introduced two simplifications to fit the connectivity and size constraints of the Xylo chip. First, by avoiding input copies and by collapsing the ON and OFF channels of each tactile sensor into a single spike array through an OR-like operation at every timestep, we reduced the input size to twelve. Second, we selected a subset of seven possible characters (’Space’, ’A’, ’E’, ’I’, ’O’, ’U’, ’Y’), thus making this a spatio-temporal classification problem with seven classes. For the recurrent layer of the spiking recurrent neural network (SRNN), 38 (zero, bias) or 40 (subtractive, no bias) CuBa-LIF neurons were used. To identify optimal hyperparameters for the SRNN, an optimization procedure adapted from the one described in [49] was performed. The network was trained with backpropagation-through-time using surrogate gradients in snnTorch. To optimize the spiking activity of the output neurons, training was performed on the cross entropy of the spike count at the output layer. The spike function was implemented with the Heaviside function in the forward pass and approximated with the fast sigmoid function in the backward pass [50]: $S\approx\frac{U}{1+k|U|}$ with $\frac{\partial S}{\partial U}=\frac{1}{(1+k|U|)^{2}}$ where $k=5$. A regularization term was added to this loss function to penalize both the $\ell^{1}$ norm of the average number of spikes per neuron and the $\ell^{2}$ norm of the total number of spikes in the recurrent layer. The coefficients for the regularization terms were chosen through the above-mentioned hyperparameter optimization, and the optimal values found were $\mu_{L1}=6\times 10^{-4}$, $\mu_{L2}=4\times 10^{-6}$ (zero, bias) and $\mu_{L1}=1\times 10^{-3}$, $\mu_{L2}=1\times 10^{-6}$ (subtractive, no bias). The objective was minimized using the Adam optimizer with a learning rate of 0.005 (zero, bias) or 0.001 (subtractive, no bias). The training proceeded for 500 epochs, and the parameters that yielded the highest validation accuracy were chosen.
5 Acknowledgements
The authors would like to acknowledge the Telluride Neuromorphic Cognition Engineering Workshop for providing a conducive environment for conceiving and implementing the first steps towards NIR. J.E.P. acknowledges funding from the EC Horizon 2020 Framework Programme under Grant Agreements 785907 and 945539 (HBP) as well as the Pioneer Centre for AI, under the Danish National Research Foundation grant number P1. S.A. acknowledges funding from the European Union’s Horizon 2020 Research and Innovation Programme under the Marie Skłodowska-Curie Grant Agreement No. 860360 (POST DIGITAL). M.J. acknowledges funding by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) as part of Germany’s Excellence Strategy – EXC 2050/1 – Project ID 390696704 – Cluster of Excellence ”Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of Technische Universität Dresden. BV acknowledges funding by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) under the contract 01MN23004F (ESCADE). V.F. acknowledges funding by the 3A-ITALY project (CUP E13C22001900001). G.U. acknowledges funding by the Ebrains-Italy project (CUP B51E22000150006). J.K.E. acknowledges funding from the National Science Foundation under Award ECCS-2332166. D.R.M and S.S acknowledge the ECSEL Joint Undertaking grant agreements 826655 “TEMPO” and 876925 “ANDANTE”; the KDT Joint Undertaking grant agreement 101097300 “EdgeAI”; Innosuisse and the Swiss State Secretariat for Education, Research and Innovation (SERI).
6 Contributions
J.E.P., S.A., M.J., F.C.B., B.V., G.L. and S.S. contributed to the NIR Core library. J.E.P., S.A., M.J., V.F., F.C.B., B.V., G.L., D.R.M., P.Z., K.H., T.C.S., J.K.E. and S.S. contributed to the conversions and experiments for the different library and hardware backends. J.E.P., S.A., B.V., S.S., G.U. and V.F. conceived the experiments. J.E.P., S.A., M.J., and J.K.E. contributed to the figures. All authors contributed to writing the paper.
7 Competing interests
The authors declare no competing interests.
8 Data availability
All data and code used in this paper is available at https://github.com/neuromorphs/nir.